# KCL-PH-TH/2025-33

### v1

@@ III.1 Sub-Horizon and Near-Horizon Scale Perturbations
In this section, we investigate the effects of both large and small amplitude inhomogeneities in the sub-horizon regime, where $\lambda<H_{0}^{-1}$. In the perturbation theory limit, sub-horizon modes evolve as $\rho_{\delta\phi}\sim a^{-4}$, indicating a behaviour that resembles radiation. According to [23], this behaviour is robust regardless of whether the inhomogeneities are in the linear or non-linear regime in the presence of a fixed background. We aim to investigate the behaviour of the scalar self-perturbation in a pure kination regime, allowing the inhomogeneities to evolve into the non-linear regime and potentially become the dominant energy density contribution. As stated above, one normally expects the presence of a subdominant potential term, which may nevertheless remain subdominant in the post-inflationary evolution compared to the kinetic energy density and can therefore be safely neglected.
@@
We simulate a scalar field with an initially homogeneous kinetic energy profile, superimposed with an inhomogeneity characterized by an initial wavelength of $\lambda=0.01H_{0}^{-1}$ as in Eq. (6). The amplitude of this inhomogeneity is chosen such that the volume-averaged initial gradient energy density is a tenth of the kinetic energy density, i.e. $b=0.1$. To focus on the localised dynamics, we capture only a single full wavelength of the perturbation within our simulation domain instead of the entire Hubble volume. We study the evolution of this perturbation by tracking the volume-averaged perturbation energy density defined in Eq. (21). We present our results in Figure 1.
@@
Figure 1: Simulation results for $\lambda_{0}=0.01H_{0}^{-1}$, $b=0.1$ – at $a\sim 3$, the “perturbation” energy $\rho_{\delta\phi}$ dominates over the “background” energy density, but the two “components” still evolve as $a^{-4}$ and $a^{-6}$ respectively. Upper panel: the evolution of volume-averaged energy densities of the background and perturbation. Lower panel: background and perturbation energy densities as a power law of the volume-averaged scale factor $a=\langle\chi\rangle^{-1/2}$.
@@
We found that a sub-horizon mode in the subdominant regime indeed behaves as $\langle\rho_{\delta\phi}\rangle\sim a^{-4}$. It redshifts as $a^{-4}$, becomes ($)\order{1}$ with the kination background, and dominates as an effective radiation-like background. The continuity in the energy density profiles in Figure 1 shows that the transition from kination domination to radiation domination is smooth.
@@
As [38, 39] has shown that direct collapse black hole formation is possible with a supercritical sub-horizon initial overdensity during matter domination, we investigate such processes for kination. We discovered that for inhomogeneities with characteristic wavelengths of $\lambda\sim\order{10^{-1}}H_{0}^{-1}$ the critical overdensity is of ($)\order{10^{2}\sim 10^{3}}$. In such conditions, the universe’s evolution is dominated by a highly localised gradient energy density of a massless scalar field. Therefore, gravitational collapse via accretion is highly suppressed. We thus conclude that, even if sub-horizon perturbations manage to grow during a kination era to become $\mathcal{O}(1)$, it is still very unlikely for such perturbations to undergo gravitational collapse.
@@
As we probe sub-horizon inhomogeneities with wavelengths closer to the horizon size, the critical value $b_{\text{crit}}$ for black hole formation is reduced. For $\lambda_{0}=0.5H_{0}^{-1}$, we find $b_{\text{crit}}\approx 20$, which still implies that gradients must initially be very dominant. Once black holes form in our simulation, the universe becomes matter-dominated, with black holes forming on a lattice due to the symmetries enforced by the periodic boundary conditions.
@@
In the near-Hubble limit, where $\lambda_{0}=H_{0}^{-1}$, we find that black hole formation can occur more easily, with a critical value $b_{\text{crit}}\approx 2$. We note that while the wavelength is nominally of Hubble size, the significant amount of energy density from the gradient of the scalar field effectively shrinks the actual Hubble sphere, rendering the inhomogeneity effectively super-horizon. In the limit of such large inhomogeneities at this length scale, it is not meaningful to associate the energy density evolution with the scale factor, nor is the scale factor itself well-defined due to the lack of homogeneity. In cases where black holes do not form, we observe that the energy density eventually, on average, evolves as in a radiation-dominated universe. Additionally, for near-Hubble wavelengths, with $b\ll 1$, the perturbations grow relative to the kination background and eventually dominate the evolution, again behaving like a radiation-dominated universe in the same manner as in the deep sub-horizon case.
@@
To conclude this section: in the limit where a kination background with small perturbations is present, the perturbation theory result remains valid—namely, small perturbations redshift more slowly than the kinetic background and eventually dominate the dynamics of the universe, becoming a new ”background” that behaves like radiation, with the kinetic background becoming subdominant. On the other hand, when the initial $b\gtrsim 1$, inhomogeneities dominate the evolution, but the gravitational backreaction is not strong enough to induce large-scale collapse. However, we checked that in the limit $b\gg 1$, gravitational collapse can occur although this is not a surprise : sufficiently large homogeneities will always dominate local dynamics and cause collapse.
III.2 Super-Horizon Perturbations
@@
For inhomogeneities with characteristic wavelength $\lambda_{0}>H_{0}^{-1}$, perturbation theory predicts that the perturbation energy density of super-horizon modes is approximately given by the gradient term and scales as $\rho_{\delta\phi}\sim\rho_{\nabla}\sim a^{-2}$. Thus, it effectively “grows” as $a^{4}$ compared to the kination background and behaves like scalar curvature perturbations. However, in practice, as shown in the previous section that the energy density of the perturbations consists of both the gradient term and the induced kinetic term. Thus, we aim to investigate whether the perturbation energy density and/or the gradient energy density still behave as $\rho_{\delta\phi}\sim a^{-2}$ in the presence of induced kinetic inhomogeneity and strong gravitational backreaction.
@@
We begin by confirming this in the linear limit and simulate a scalar field with a homogeneous kinetic energy profile and an inhomogeneity characterised by an initial wavelength of $\lambda_{0}=10^{4}H_{0}^{-1}$. The amplitude of this inhomogeneity is chosen by setting the non-linearity parameter to $b=10^{-10}$. The results of the simulation are shown in Figure 2. We find that once the system relaxes from the initial conditions, the volume-averaged perturbation energy density closely tracks the gradient energy density. We measure the variance of the local expansion parameter $\langle\delta K\rangle/\langle K\rangle\ll 1$, indicating that the simulation domain remains highly homogeneous and the spacetime closely follows the FLRW metric. We are thus justified in identifying $\langle\chi\rangle^{-1/2}\sim a$ and the perturbation energy density does evolve as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$ in this linear regime.
@@
Figure 2: Simulation results for $\lambda_{0}=10^{4}H_{0}^{-1}$, $b=10^{-10}$. The perturbation energy density is closely tracked by the gradient energy density and evolve as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$ while the kination background remains at $\langle\overline{\rho}\rangle\sim a^{-6}$. Upper panel: the evolution of volume-averaged energy densities of the background, the gradient, and perturbations, Lower panel: background, gradient, and perturbation energy densities as a power law of the volume-averaged scale factor $a=\langle\chi\rangle^{-1/2}$.
@@
We extend the simulation duration to investigate whether these perturbations continue to behave as $\langle\rho_{\delta\phi}\rangle\sim a^{-2}$ when they enter a non-linear regime. Since the oscillation period of the super-horizon mode is proportional to the characteristic wavelength, $T\propto\lambda\gg H^{-1}$, for the mode to oscillate and eventually re-enter the horizon to potentially enter a strong-gravity regime, we shorten the initial characteristic wavelength and increase the initial gradient energy density. We present a representative case with $\lambda_{0}=200H_{0}^{-1}$ and $b=1.8\times 10^{-6}$ in Figure 3.
@@
Figure 3: Simulation results for $\lambda_{0}=200H_{0}^{-1}$, $b=1.8\times 10^{-6}$. The formation of black hole is indicated by the discontinuity around $a_{\mathrm{BH}}\sim 22$ at the end of the plot. The perturbation energy density deviates away from the perturbative description and shifts towards $\langle\rho_{\delta\phi}\rangle\sim a^{-1}$, with the gradient energy density fluctuates around $\langle\rho_{\nabla}\rangle\sim a^{-2}$. Upper panel: the evolution of volume-averaged energy densities of the background, gradient, kinetic, and perturbations, Lower panel: The background, gradient, kinetic, and perturbation energy densities as a power law of the volume-averaged conformal factor $\langle\chi\rangle^{-1/2}$. Simulation movie: https://youtu.be/rPnIBkoagwM
@@
We observe that while the perturbation energy density is initially “attracted” to $\langle\rho_{\delta\phi}\rangle\sim a^{-2}$ initially as perturbation theory suggests, the behaviours of both $\langle\rho_{\delta\phi}\rangle$ and $\langle\rho_{\nabla}\rangle$ quickly deviate from a constant power law once the oscillatory behaviour of the super-horizon mode and the induced kinetic perturbation energy density becomes significant. In particular, the behaviour for the total perturbation evolves approximately as $\langle\rho_{\delta\phi}\rangle\sim a^{-1}$, while the gradient term exhibits an oscillatory behaviour around $\langle\rho_{\nabla}\rangle\sim a^{-2}$. This indicates that the oscillatory behaviour and strong gravitational backreaction are significant.
@@
$a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}\,.$ (26)
This estimate assumes a consistent and well-defined notion of the scale factor throughout the simulation domain and evolution. To ensure this, we verify that the domain remains well-described by an FLRW spacetime. Following the methods of [40, 36] we evaluate the variance, $\sigma^{2}$, of the local expansion parameter. In scenarios where black hole formation is expected, we additionally monitor the variance of the conformal factor and compute their kurtoses, $g_{2}$ [^2] We use the natural estimator definition of $g_{2}=\frac{\frac{1}{n}\sum_{i}^{n}(x_{i}-\bar{x})^{4}}{\sigma^{4}}-3,$ where $\sigma$ is the standard deviation. Under this definition, $g_{2}=-3$ indicates absolute uniformity, $g_{2}=0$ indicates a normal distribution, and $g_{2}>0$ indicates a sharp peak in distribution., to assess deviation from FLRW behaviour near the onset of gravitational collapse. In the simulation shown in Figure 3, we observe that both $\sigma^{2}_{K}$ and $\sigma^{2}_{\chi}$ remain in the range ($)\order{10^{-3}\sim 10^{-5}}$ throughout the evolution. We also find that $g_{2,\chi}\sim g_{2,K}\sim-1$ prior to the black hole formation showing platykurtic(or near uniformity) distribution. These results confirm that the spacetime remains an excellent approximation to FLRW before the onset of black hole formation.
In the example shown in Figure 3, we estimate the horizon re-entry to occur at $a_{\mathrm{re-entry}}\approx 14$ with $\langle\rho_{\delta\phi}\rangle\lesssim\langle\overline{\rho}\rangle$ and black hole formation at $a_{\mathrm{BH}}\approx 22$ as indicated by the discontinuity in the energy density profiles. The first detection of an apparent horizon and the estimated mode horizon re-entry are separated by approximately 0.5 e-folds. We observed that this behaviour is consistent across all super-horizon modes we have tested.
As mentioned above, we aim to investigate whether the critical overdensity is robust for super-horizon modes that exhibit non-linear behaviour before horizon re-entry. We simulated a range of super-horizon modes with initial wavelengths spanning from $5H_{0}^{-1}$ to $2000H_{0}^{-1}$ and a non-linearity parameter $b$ ranging from ($)\order{10^{-2}\sim 10^{-9}}$. We present the results in Figure 4.
@@
Given the relative growth of the perturbation, the simulation domain deviates from pure FLRW spacetime, making the precise moment of horizon re-entry difficult to define. We define $\delta_{c,\text{ numerical}}$ as the effective critical overdensity, extrapolated from the simulation results under the assumptions that the kinetic energy density always dominates the universe’s evolution prior to horizon re-entry, and that the perturbation grows relative to the kination background remains as $a^{4}$
$\delta_{c,\text{ numerical}}\approx a^{4}b=(\lambda_{0}H_{0})^{2}b$ (27)
Therefore the presented values do not represent the critical overdensity for primordial black hole formation at horizon re-entry. Instead, they serve as a comparison to the theoretical predictions if one expects the super-horizon modes to stay frozen as in the perturbative limit, until re-entry, i.e $\delta\propto a^{4}$. Under this assumption, these are the values that correspond to the critical initial conditions that lead to primordial black holes. We detail the cosmological implications in the following section.
Figure 5: The mass of the black holes from simulations at the point of formation and a power law fit $M_{BH}\approx 0.001\times(\lambda_{0}H_{0})^{2.3}$
In the simulation where black holes form, the lightest black holes at the point of formation are shown in Figure 5 with respect to the initial Hubble scale. To compare to the Hubble scale at the point of formation, we note that while the volume-averaged perturbation energy density is of ($)\order{1}$ compared to the background, the inhomogeneities are concentrated in the proximity of the primordial black hole both before and after formation. Therefore, we are still able to extract the Hubble parameter from the homogeneous bulk region. This corresponds roughly to ($\order{10^{-2}\sim 10^{-3}}H^{-1}m_{P}^{2}$.
@@
    As discussed above, we observe the point of horizon re-entry occurs at $a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}$ with black hole formation typically taking place about 0.5$\sim$ 0.5∼ 0.5 e-folds later. In this work, the largest scale factor we were able to probe is $a\sim 80$. To investigate longer super-horizon modes that would re-enter the horizon during a prolonged kination era lasting up to $N_{\text{kin}}\sim 10$, the simulation would need to reach scale factors of ($)\order{10^{3}\sim 10^{4}}$. However, achieving such large scale factors is currently infeasible with the numerical setup employed in this study. Numerical stability constrains the maximum allowed time step, and reaching significantly larger values of $a$ leads to an accumulation of numerical errors and requires computational resources that grow exponentially with simulation duration.
@@
    The gauge choice we have imposed in Eq. (17) ensures that the presence of a small perturbation does not induce strong gravitational backreaction until close to horizon re-entry. Thus the simulation domain is effectively a uniformly expanding spacetime with a decaying expansion rate. In such domains, we observe the local expansion parameter is $K\lesssim 0$, with regions fluctuating slightly above and below the average. While Eq. (17) is a more cosmology-friendly gauge [36], the prolonged oscillation periods for longer super-horizon modes gradually drive the lapse value across the entire hyperslice downward. As detailed in II.3, the lapse function is initially set to 1 everywhere on the hyperslice. The combined effect of the dynamics and this gauge choice causes the lapse function to drop to ($)\order{10^{-1}}$ across the domain, thereby increasing both the simulation time and computational cost.
IV Implications for Cosmology

### v2

@@ 3.1 Sub-Horizon and Near-Horizon Scale Perturbations
In this section, we investigate the effects of both large and small amplitude inhomogeneities in the sub-horizon regime, where $\lambda<H_{0}^{-1}$. In the perturbative limit, sub-horizon modes evolve as $\rho_{\delta\phi}\sim a^{-4}$, indicating a behaviour that resembles radiation. According to [eröncel2025universalbounddurationkination], this behaviour is robust regardless of whether the inhomogeneities are in the linear or non-linear regime in the presence of a fixed background. As stated above, one normally expects the presence of a subdominant potential term, which may nevertheless remain subdominant in the post-inflationary evolution compared to the kinetic energy density and can therefore be safely neglected.
@@
We simulate a scalar field with an initially homogeneous kinetic energy profile, superimposed with an inhomogeneity characterised by an initial wavelength of $\lambda=0.01H_{0}^{-1}$ as in Eq. (2.6). The amplitude of this inhomogeneity is chosen such that the volume-averaged initial gradient energy density is one tenth of the background kinetic energy density, i.e. $\delta_{0}=0.1$. To focus on the localised dynamics, we capture only a single full wavelength of the perturbation within our simulation domain and apply periodic boundary condition instead of the entire Hubble volume. We present our results in Figure 1.
@@
Figure 1: Simulation results with initial gradient and kinetic perturbation in red and blue, respectively, for $\lambda_{0}=0.01H_{0}^{-1}$, $\delta_{0}=0.1$ – at $N\sim 1.2$, the “perturbation” energy $\rho_{\delta\phi}$ dominates over the “background” energy density, but the two “components” still evolve as $a^{-4}$ and $a^{-6}$ respectively. The source of the initial perturbation energy density is indistinguishable to the system as long as they are of sub-horizon. Upper panel: the evolution of volume-averaged energy densities of the background and perturbation. Lower panel: background and perturbation energy densities as a power law of the volume-averaged scale factor $a\approx\langle\chi\rangle^{-1/2}$.
@@
We found that a sub-horizon mode indeed behaves as $\langle\rho_{\delta\phi}\rangle\sim a^{-4}$, and would become $\mathcal{O}(1)$ with the kination background, and dominates as an effective radiation-like background. The continuity in the energy density profiles in Figure 1 indicates that the transition from kination domination to radiation domination is smooth.
@@
As [21, 20] has shown that critical collapse to a primordial black hole is possible with a supercritical sub-horizon initial overdensity during matter domination, we investigate such processes for kination. We discovered that for inhomogeneities with characteristic wavelengths of $\lambda\sim\mathcal{O}(10^{-1})H_{0}^{-1}$ the critical overdensity is of $\mathcal{O}(10^{2}\sim 10^{3})$. In such conditions, the universe’s evolution is dominated by a highly localised gradient energy density of a massless scalar field. Therefore, gravitational collapse via accretion is highly suppressed. We thus conclude that, even if sub-horizon perturbations manage to grow during a kination era to become $\mathcal{O}(1)$, it is still very unlikely for such perturbations to undergo gravitational collapse.
@@
As we probe sub-horizon inhomogeneities with wavelengths closer to the horizon size, the critical value $\delta_{0,\text{crit}}$ for black hole formation is reduced. For $\lambda_{0}=0.5H_{0}^{-1}$, we find $\delta_{0,\text{crit}}\approx 20$, which still implies that gradients must initially be very dominant. Once black holes form in our simulation, the universe becomes matter-dominated, with black holes forming on a lattice due to the symmetries enforced by the periodic boundary conditions.
@@
In the near-Hubble limit, where $\lambda_{0}=H_{0}^{-1}$, we find that black hole formation can occur more easily, with a critical value $\delta_{0,\text{crit}}\approx 2$. We note that while the wavelength is nominally of Hubble size, the significant amount of energy density from the gradient of the scalar field effectively shrinks the actual Hubble sphere, rendering the inhomogeneity effectively super-horizon. In the limit of such large inhomogeneities at this length scale, it is not meaningful to associate the energy density evolution with the scale factor, nor is the scale factor itself well-defined due to the lack of homogeneity. In cases where black holes do not form, we observe that the energy density eventually, on average, evolves as in a radiation-dominated universe. Additionally, for near-Hubble wavelengths, with $\delta_{0}<1$, the perturbations grow relative to the kination background and eventually dominate the evolution, again behaving like a radiation-dominated universe in the same manner as in the deep sub-horizon case.
@@
To conclude this section: in the limit where a kination background with small perturbations is present, the perturbation theory result remains valid. On the other hand, when the initial $\delta_{0}\gtrsim 1$, inhomogeneities dominate the evolution, but the gravitational backreaction is not strong enough to induce large-scale collapse. Finally, in the limit $\delta_{0}\gg 1$, gravitational collapse can occur as sufficiently large inhomogeneities will always dominate local dynamics.
3.2 Super-Horizon Gradient Perturbations
@@
For inhomogeneities with characteristic wavelength $\lambda_{0}>H_{0}^{-1}$, linear perturbation theory indicates that the perturbation energy density of super-horizon modes is approximately given by the gradient term and scales as $\rho_{\delta\phi}\sim\rho_{\nabla}\sim a^{-2}$. Thus, the density contrast grows as $a^{4}$ compared to the kination background and the modes behave like scalar curvature perturbations. However, in practice, as shown in the previous section, the energy density of the perturbations consists of both the gradient term and the induced kinetic term, but the superhorizon kinetic part according to the perturbative analysis tracks the background and decays as $\rho_{\delta K}\sim a^{-6}$, therefore, very quickly dilutes, we
@@
We begin by confirming this in the linear limit and simulate a scalar field with a homogeneous kinetic energy profile and an inhomogeneity characterised by an initial wavelength of $\lambda_{0}=2\times 10^{4}H_{0}^{-1}$. The amplitude of this inhomogeneity is chosen by setting the initial density contrast to $\delta_{0}=4\times 10^{-10}$. The results of the simulation are shown in Figure 2. We find that the volume-averaged perturbation energy density $\langle\rho_{\delta\phi}\rangle$ very quickly tracks the gradient energy density once the system relaxes from the initial conditions. This happens within approximately one e-fold. We also find that the variance of the local expansion parameter $\langle\delta K\rangle/\langle K\rangle\ll 1$, indicating that the simulation domain remains highly homogeneous and the spacetime closely follows the FLRW metric. Thus, the perturbation energy density then evolves as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$.
@@
Figure 2: Simulation results for $\lambda_{0}=1\times 10^{3}H_{0}^{-1}$, $\delta_{0}=1\times 10^{-10}$. The perturbation energy density is closely tracked by the gradient energy density and evolve as $\langle\rho_{\delta\phi}\rangle\sim\langle\rho_{\nabla}\rangle\sim a^{-2}$ while the kination background remains at $\overline{\rho}\sim a^{-6}$. Upper panel: the evolution of volume-averaged energy densities of the background, the gradient, and perturbations, Lower panel: background, gradient, and perturbation energy densities as a power law of the volume-averaged scale factor $a\approx\langle\chi\rangle^{-1/2}$.
@@
Since the oscillation period of a super-horizon mode is proportional to its characteristic wavelength, $T\propto\lambda\gg H^{-1}$, the mode must oscillate and eventually re-enter the horizon before it can potentially enter a strong-gravity regime. To facilitate this within a feasible simulation time, we reduce the initial characteristic wavelength and increase the initial gradient energy density. As a representative example, we consider the case $\lambda_{0}=20H_{0}^{-1}$ and $\delta_{0}=5\times 10^{-4}$, shown in Figure 3.
@@
Figure 3: Simulation results for $\lambda_{0}=20H_{0}^{-1}$, $\delta_{0}=5\times 10^{-4}$. The formation of black hole is indicated by the discontinuity around $\log(a)\sim 1.9$ at the end of the plot. The perturbation energy density deviates upwards away from the perturbative description when the induced kinetic perturbation becomes comparable with the gradient perturbation energy density. Upper panel: the evolution of volume-averaged energy densities of the background, gradient, kinetic, and perturbations, Lower panel: The background, gradient, kinetic, and perturbation energy densities as a power law of the $\log$ of the volume-averaged scale factor $a\approx\langle\chi\rangle^{-1/2}$. Simulation movie: https://www.youtube.com/watch?v=0msEzENelt8
@@
We observe that $\langle\rho_{\delta\phi}\rangle\sim a^{-2}$ initially, the behaviours of both $\langle\rho_{\delta\phi}\rangle$ and $\langle\rho_{\nabla}\rangle$ quickly deviate from a constant power law once the oscillatory behaviour of the super-horizon mode and the induced kinetic perturbation energy density becomes significant. In particular, the behaviour for the total perturbation evolves approximately as $\langle\rho_{\delta\phi}\rangle\sim a^{-1}$, while the gradient term exhibits an oscillatory behaviour around $\langle\rho_{\nabla}\rangle\sim a^{-2}$. This indicates that the oscillatory behaviour and strong gravitational backreaction are significant.
@@
$a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}\,.$ (3.1)
This estimate assumes a consistent and well-defined notion of the scale factor throughout the simulation domain and evolution. To ensure this, we verify that the domain remains well-described by an FLRW spacetime. Following the methods of [27, 7] we evaluate the variance, $\sigma^{2}$, of the local expansion parameter. In scenarios where black hole formation is expected, we additionally monitor the variance of the conformal factor and compute their kurtoses, $g_{2}$ [^2] We use the natural estimator definition of $g_{2}=\frac{\frac{1}{n}\sum_{i}^{n}(x_{i}-\bar{x})^{4}}{\sigma^{4}}-3,$ where $\sigma$ is the standard deviation. Under this definition, $g_{2}=-3$ indicates absolute uniformity, $g_{2}=0$ indicates a normal distribution, and $g_{2}>0$ indicates a sharp peak in distribution., to assess deviation from FLRW behaviour near the onset of gravitational collapse. In the simulation shown in Figure 3, we observe that both $\sigma^{2}_{K}$ and $\sigma^{2}_{\chi}$ remain in the range $\mathcal{O}(10^{-3}\sim 10^{-5})$ throughout the evolution. We also find that $g_{2,\chi}\sim g_{2,K}\sim-1$ prior to the black hole formation showing platykurtic(or near uniformity) distribution. These results confirm that the spacetime remains a good approximation to FLRW prior to the black hole formation.
In the simulation example shown in Figure 3, we estimate the horizon re-entry to occur at $\log(a)_{\mathrm{re-entry}}\approx 1.5$ with $\langle\rho_{\delta\phi}\rangle\lesssim\overline{\rho}$ and black hole formation at $\log(a)_{\mathrm{BH}}\approx 1.9$ as indicated by the discontinuity in the energy density profiles. The first detection of an apparent horizon and the estimated mode horizon re-entry are separated by approximately 0.5 e-folds. We observed that this behaviour is consistent across all super-horizon modes collapses we have tested.
Given the relative growth of the perturbation, the simulation domain deviates from pure FLRW spacetime, making the precise moment of horizon re-entry difficult to define. We define $\delta_{c,\text{ effective}}$ as the effective critical overdensity, extrapolated from the simulation results under the assumptions that the kinetic energy density always dominates the universe’s evolution prior to horizon re-entry, and that the perturbation grows relative to the kination background remains as $a^{4}$
$\delta_{c,\text{ effective}}\approx a^{4}\delta_{0,\text{crit}}=(\lambda_{0}H_{0})^{2}\delta_{0,\text{crit}}$ (3.2)
Note that this quantity is defined for the sake of comparing our results to the perturbative case, but does not correspond to a quantity that is actually measurable numerically, unlike $\delta_{0,\text{crit}}$, which is the physical and measurable quantity in our approach. As mentioned above, we aim to investigate whether the critical overdensity is robust for super-horizon modes that exhibit non-linear behaviour before horizon re-entry. We simulated a range of super-horizon modes with initial wavelengths spanning from $5H_{0}^{-1}$ to $2000H_{0}^{-1}$ and an initial density contrast $\delta_{0}$ ranging from $\mathcal{O}(10^{-2}\sim 10^{-9})$. We present the results in Figure 4.
@@
Note that the presented values do not represent the critical overdensity for primordial black hole formation at horizon re-entry. Instead, they serve as a comparison to the theoretical predictions if one expects the super-horizon modes to stay frozen as in the perturbative limit, until re-entry, i.e $\delta\sim a^{4}$. Under this assumption, these are the values that correspond to the critical initial conditions that lead to primordial black holes. We detail the cosmological implications in the following section.
Figure 5: The mass of the black holes from simulations at the time of formation and a power law fit $M_{BH}\approx 0.001\times(\lambda_{0}H_{0})^{2.3}$
In the simulation where black holes form, the lightest black holes at the point of formation are shown in Figure 5 with respect to the initial Hubble scale. To compare to the Hubble scale at the point of formation, we note that while the volume-averaged perturbation energy density is of $\mathcal{O}(1)$ compared to the background, the inhomogeneities are concentrated in the proximity of the primordial black hole both before and after formation. Therefore, we are still able to extract the Hubble parameter from the homogeneous bulk region. This corresponds roughly to $\mathcal{O}(10^{-2}\sim 10^{-3})H^{-1}m_{P}^{2}$.
@@
    As discussed above, we observe the point of horizon re-entry occurs at $a_{\text{re-entry}}\approx\sqrt{\lambda_{0}H_{0}}$ with black hole formation typically taking place about $\sim 0.5$ e-folds later. In this work, the largest scale factor we were able to probe is $a\sim 80$. To investigate longer super-horizon modes that would re-enter the horizon during a prolonged kination era lasting up to $N_{\text{kin}}\sim 10$, the simulation would need to reach scale factors of $\mathcal{O}(10^{3}\sim 10^{4})$. However, achieving such large scale factors is currently infeasible with the numerical setup employed in this study. Numerical stability constrains the maximum allowed time step, and reaching significantly larger values of $a$ leads to an accumulation of numerical errors and requires computational resources that grow exponentially with simulation duration.
@@
    The gauge choice we have imposed in Eq. (2.17) ensures that the presence of a small perturbation does not induce strong gravitational backreaction until close to horizon re-entry. Thus the simulation domain is effectively a uniformly expanding spacetime with a decaying expansion rate. In such domains, we observe the local expansion parameter is $K\lesssim 0$, with regions fluctuating slightly above and below the average. While Eq. (2.17) is a more cosmology-friendly gauge [7], the prolonged oscillation periods for longer super-horizon modes gradually drive the lapse value across the entire hyperslice downward. As detailed in 2.3, the lapse function is initially set to 1 everywhere on the hyperslice. The combined effect of the dynamics and this gauge choice causes the lapse function to drop to $\mathcal{O}(10^{-1})$ across the domain, thereby increasing both the simulation time and computational cost.
3.3 Super-Horizon Kinetic Perturbations
For completeness, we also investigated the effects of a pure kinetic super-horizon perturbation in this regime. For such a mode, perturbation theory suggests $\rho_{\delta K}\sim a^{-6}$ (see Appendix. A). However, as we have seen in the previous section that the super-horizon gradient perturbation redshifts as $a^{-2}$, the induced super-horizon gradient mode would eventually become the dominant contribution of the scalar perturbation. We present the result from the simulation with the same $\lambda_{0}$ and $\delta_{0}$ as in Figure 3. We observe that initially, the kinetic perturbation drops drastically to induce gradient energy until they are of the same order. The total perturbation energy density then behaves roughly as $a^{-2}$ before the estimated re-entry point. However, since the gradient energy density is induced as a later time than in the simulation of Figure 3 and at a lower order, it does not have enough time to grow to the same non-linear regime as in the previous simulation, and thus do not undergo gravitational collapse. After the estimated re-entry at around $\log(a)\sim 1.5$, it redshifts as $a^{-4}$, which again confirms the result for the sub-horizon case.
Figure 6: Simulation results for super-horizon kinetic perturbation with initial conditions of $\lambda_{0}=20H_{0}^{-1}$, $\delta_{0}=5\times 10^{-4}$. Upper panel: the evolution of volume-averaged energy densities of the background, gradient, kinetic, and perturbations, Lower panel: The background, gradient, kinetic, and perturbation energy densities as a power law of the $\log$ of the volume-averaged scale factor $a=\langle\chi\rangle^{-1/2}$.
4 Implications for Cosmology